OpenStudy (anonymous):

what is the common ratio of a sequence 2, -6, 18,-54,...?

OpenStudy (anonymous):

-3

OpenStudy (whpalmer4):

To find the common ratio, just divide two adjacent elements. \[\frac{-6}{2} =\]\[\frac{18}{-6}=\]\[\frac{-54}{18}=\]See a pattern? :-)